# Rotation sets and axes in the fine curve graph for torus homeomorphisms

### v1

@@ 3.4. Bestvina-Fujiwara equivalence relation
Let $G$ be a group acting on a $\delta$-hyperbolic space, and $f,g\in G$. Assume they act loxodromically, and denote $\xi^{\pm}(f),\xi^{\pm}(g)$ their end-points. Bestvina-Fujiwara ([BF02]) define a relation $f\sim g$ if there are copies of the axis of $g$ arbitrarily close to the axis of $f$. More precisely: if for every neighborhoods $V^{-},V^{+}$ of $\xi^{-}(f),\xi^{+}(f)$ there exists $h\in G$ such that $h(\xi^{-}(g))\in V^{-},h(\xi^{+}(g))\in V^{+}$.
@@ 4. Fine curve graphs
The key point is that bicorn surgery paths are geometrically efficient. This follows from work of Rasmussen [Ras20] on nonseparating curve graphs; see [BHW24, Section 4] for details.
Lemma 4.2 ([BHW24, Lemma 4.2]).
@@
We will also make use of the fine arc graph of the annulus $A=[0,1]\times S^{1}$, denoted $\mathcal{A}^{\dagger}(A)$, whose vertices are essential proper arcs in $A$, that is, images of $[0,1]$ under an embedding that maps $(0,1)$ into the interior of $A$ and 0 and $1$ to distinct boundary components of $A$, and edges correspond to disjointness. The distance between two vertices $a\neq b$ in this graph may be computed as follows (see [BHM+22, LRW24]).
@@ 5. Fine markings
Let $m_{0}=(\alpha_{0},\beta_{0})$. We choose a covering of the surface by disks $U_{i}$ which are: the complement $U_{1}$ of $\alpha_{0}\cup\beta_{0}$, a small neighbourhood $U_{2}$ of $\alpha_{0}\cap\beta_{0}$, and neighbourhoods $U_{3},U_{4}$ of the arcs $\alpha_{0}\cup\beta_{0}\setminus U_{2}$. By the work of Mann-Rosendal ([MR18]), the geometry of $\mathrm{Homeo}_{0}(T^{2})$ is quasi-isometric to the fragmentation norm induced by the cover $\{U_{i}\}$. We denote this norm by $\|\cdot\|$. We denote the metric on $\mathcal{M}^{\dagger}(m_{0})$ for brevity by $d$. By right-invariance of the metric on $\mathrm{Homeo}_{0}(T^{2})$ and of the action by inverses of maps on $\mathcal{M}^{\dagger}(m_{0})$, it suffices to estimate $d(m_{0},\phi^{-1}m_{0})$ in terms of $\|\phi\|$, which we will do now.
@@
Figure 7. Moving in the marking graph. The lines between consecutive $f_{1}\cdots f_{i}m_{0},f_{1}\cdots f_{i}m_{0}$ indicate distance at most $4$.
Now, on the other hand, suppose that $m_{1},m_{0}$ differ by a single elementary move. Then there is an element $f\in\mathrm{Homeo}_{0}(T^{2})$ with $f(m_{0})=m_{1}$ and $\|f\|\leq 4$; in fact $f$ can be written as a product of homeomorphisms supported on each $U_{i}$ at most once. Namely, assume that $\alpha_{0}$ is unchanged and $\beta_{0}$ changes to $\beta_{1}$, with $\beta_{0}\cap\beta_{1}=\emptyset$. By a modification in the neighbourhood of the intersection point followed by another one in the neighourhood of $\beta_{0}$, we get a curve $f_{1}f_{0}(\beta_{0})$ which is disjoint from $\beta_{0}$. By applying a homeomorphism $f_{2}$ supported in the neighborhood of $\alpha_{0}$, we get $f_{2}f_{1}f_{0}(\beta_{0})$ which additionaly coincides with $\beta_{1}$ near $\alpha_{0}$. Now a modification on the complement of $\alpha_{0}\cup\beta_{0}$ (i.e. a homeomorphism $f_{3}$ of fragmentation norm $1$) can send $f_{2}f_{1}f_{0}(\beta_{0})$ to $\beta_{1}$. The existence of $f_{2}$ and $f_{3}$ is an application of a Schoenflies-like theorem (see for instance [Sie05] for an account of the classical case, and [Hom53] for a very general case.
@@
We note that this corollary proves the easy direction of the following result due to Emmanuel Militon [Mil14]. Here $\mathrm{diam}$ denotes the diameter in the plane.
Theorem 5.4 ([Mil14, Theorem 2.4]).
@@
The phenomenon we exhibited here is completely analogous to behavior seen in mapping class groups. Here, Broaddus–Farb–Putman show that the kernel of the Birman exact sequence is exponentially distorted [BFP11].
@@ 6. Fine twist numbers and bounded geodesic image theorems
In the beginning of this section and the next one we discuss a general (orientable) surface $S$ of genus $\geq 1$. We will need only the case of the torus, but the tools developed here are likely useful in the general setting. In higher genus, edges of the curve graph corresponds to disjointness, see [BHW22].
@@ 6.2. Bounded geodesic image theorems
We have the following bounded geodesic image theorems, optimizing either for minimizing the projection bound, or the distance to the projection annulus. Both of these are inspired by Masur-Minsky’s bounded geodesic image theorem, originally proved in [MM00]. Also compare [LT25] for a bounded geodesic image theorem in higher genus fine curve graphs projecting to non-annular subsurfaces. In the following lemmas, $\mathrm{diam}$ denotes the diameter of sets of vertices in ${\mathcal{A}^{\dagger}}(A)$, and $B_{r}(\alpha)$ denotes a ball in the fine curve graph.
@@
We thank Katie Mann for pointing out to us that this second version follows easily from the first one; our initial proof was an adaptation of Richard Webb’s argument in the usual curve graph ([Web15]).
@@ 7.2. Quasi-paths
Let us first give a more concrete, geometrical characterization of being at almost distance at most $1$, which will show in particular that this is a symmetric notion. From now on we fix an orientation on our base curve $\alpha_{0}$, and orient each curve isotopic to $\alpha_{0}$ accordingly, so that the oriented curves are isotopic. Consider the infinite cyclic cover $A=\mathbb{R}\times\mathbb{S}^{1}$ as above where $\alpha_{0}$ lifts to a closed curve, and let $T:A\to A$ denote the covering transformation $(x,y)\mapsto(x+1,y)$. We may choose the orientation so that for each lift $\widetilde{\alpha}$ of $\alpha\in\mathcal{C}^{\dagger}_{0}(T^{2})$, the lift $T\widetilde{\alpha}$ is on the right-hand side of $\widetilde{\alpha}$. Given two essential simple closed curves $a,b$ in $A$, we write $a<b$ if $b$ is included in the open domain bounded by $a$ on its right-hand side, and $a\leq b$ if $b$ is included in the closure of this domain.
@@ 7.4. Operations on curves in the infinite annulus
Denote $\mathcal{C}^{\dagger}(A)$ the set of essential simple closed curve in $A$, which corresponds to curves in the plane that surrounds $\infty$. Given two curves $\alpha,\beta\in\mathcal{C}^{\dagger}(A)$, we denote $\alpha\wedge\beta$ the boundary of $R(\alpha\cup\beta)$. A theorem of Kerekjarto (see [vK23]) asserts that $\alpha\wedge\beta$ is an element of $\mathcal{C}^{\dagger}(A)$, the remarkable fact being that this holds without any regularity nor transversality assumption. Thus we have an operation $\wedge:\mathcal{C}^{\dagger}(A)\times\mathcal{C}^{\dagger}(A)\to\mathcal{C}^{\dagger}(A)$. Note the following caracterization, which is an immediate consequence of the Schoenflies theorem: a point $x$ belongs to $\alpha\wedge\beta$ if and only if (1) it belongs to $\alpha\cup\beta$, and (2) it is accessible from $\infty$ in $R(\alpha)\cap R(\beta)$, that is, there is an embedding with $\rho(0)=x$, $\rho(+\infty)=+\infty$, and $\rho((0,1))\subset R(\alpha)\cap R(\beta)$. We call $\rho$ a witness ray for $x$ in $\alpha\wedge\beta$.
@@ 8. The marking graph estimate
The aim of this section is to prove an upper bound for the distance $d(m,m^{\prime})$ in the marking graph inspired by foundational Masur-Minsky distance formula [MM00] in the “usual” marking graph.
@@
9. Metric WPD property, $C^{0}$ distance estimates
@@ 9.2. Proof of the metric WPD property
This subsection is dedicated to the proof of the metric WPD property (Proposition 1.3).
@@
In short, we get the result by showing that $h=g^{-m}f^{n}$ belongs to $Z_{2b}(\alpha,f^{n}\alpha)$, and then applying the WPD. Here are the details. Let $b,f,\alpha$ be as in the statement. By Proposition 1.3, $f$ satisfies the metric WPD property. Applying this to $\alpha$ and $b$ provides a positive integer $n_{0}$ and a positive number $D$. Let $g\in\mathrm{Homeo}_{0}(S)$, $\beta\in\mathcal{C}^{\dagger}(S)$, $n\geq n_{0}$, $m\in\mathbb{Z}$, and consider the map $h=g^{-m}f^{n}$. By the triangular inequality and the fact that $f$ and $g$ act by isometries, we deduce that $h\in Z_{2b}(\alpha,f^{n}\alpha)$: indeed,
@@ 10.3. Semi-continuity
Compared to the proofs of Theorems 3 and 4, which essentially involved only the two maps $f$ and $g$, the main difficulty here is that we need to control the quality of the quasigeodesic axes for the $g_{k}$’s. This will be achieved via Lemma 3.6.
@@ 10.4. Metric characterization of fixed points of loxodromic
The following statement contains Corollary 1.4 from the introduction.
@@
Let $D=\widetilde{d}(h,\mathrm{Id})$, and let us fix a lift $\widetilde{\alpha}_{0}$ of $\alpha_{0}$ in the universal cover. From the hypothesis it follows that $h\alpha_{0}$ has a lift which is between $\widetilde{\alpha}_{0}-(0,D)$ and $\widetilde{\alpha}_{0}+(0,D)$. This lifts meets at most $2D+1$ lifts of $\alpha_{0}$. Now the claim follows from [BHM+22, Lemma 4.5]. ∎

### v2

@@ 3.4. Bestvina-Fujiwara equivalence relation
Let $G$ be a group acting on a $\delta$-hyperbolic space, and $f,g\in G$. Assume they act loxodromically, and denote $\xi^{\pm}(f),\xi^{\pm}(g)$ their end-points. Bestvina-Fujiwara ([BF]) define a relation $f\sim g$ if there are copies of the axis of $g$ arbitrarily close to the axis of $f$. More precisely: if for every neighborhoods $V^{-},V^{+}$ of $\xi^{-}(f),\xi^{+}(f)$ there exists $h\in G$ such that $h(\xi^{-}(g))\in V^{-},h(\xi^{+}(g))\in V^{+}$.
@@ 4. Fine curve graphs
The key point is that bicorn surgery paths are geometrically efficient. This follows from work of Rasmussen [Rasmussen] on nonseparating curve graphs; see [Dagger3, Section 4] for details.
Lemma 4.2 ([Dagger3, Lemma 4.2]).
@@
We will also make use of the fine arc graph of the annulus $A=[0,1]\times S^{1}$, denoted $\mathcal{A}^{\dagger}(A)$, whose vertices are essential proper arcs in $A$, that is, images of $[0,1]$ under an embedding that maps $(0,1)$ into the interior of $A$ and 0 and $1$ to distinct boundary components of $A$, and edges correspond to disjointness. The distance between two vertices $a\neq b$ in this graph may be computed as follows (see [Dagger2, LW]).
@@ 5. Fine markings
Let $m_{0}=(\alpha_{0},\beta_{0})$. We choose a covering of the surface by disks $U_{i}$ which are: the complement $U_{1}$ of $\alpha_{0}\cup\beta_{0}$, a small neighbourhood $U_{2}$ of $\alpha_{0}\cap\beta_{0}$, and neighbourhoods $U_{3},U_{4}$ of the arcs $\alpha_{0}\cup\beta_{0}\setminus U_{2}$. By the work of Mann-Rosendal ([MR]), the geometry of $\mathrm{Homeo}_{0}(T^{2})$ is quasi-isometric to the fragmentation norm induced by the cover $\{U_{i}\}$. We denote this norm by $\|\cdot\|$. We denote the metric on $\mathcal{M}^{\dagger}(m_{0})$ for brevity by $d$. By right-invariance of the metric on $\mathrm{Homeo}_{0}(T^{2})$ and of the action by inverses of maps on $\mathcal{M}^{\dagger}(m_{0})$, it suffices to estimate $d(m_{0},\phi^{-1}m_{0})$ in terms of $\|\phi\|$, which we will do now.
@@
Figure 7. Moving in the marking graph. The lines between consecutive $f_{1}\cdots f_{i}m_{0}$ indicate distance at most $4$.
Now, on the other hand, suppose that $m_{1},m_{0}$ differ by a single elementary move. Then there is an element $f\in\mathrm{Homeo}_{0}(T^{2})$ with $f(m_{0})=m_{1}$ and $\|f\|\leq 4$; in fact $f$ can be written as a product of homeomorphisms supported on each $U_{i}$ at most once. Namely, assume that $\alpha_{0}$ is unchanged and $\beta_{0}$ changes to $\beta_{1}$, with $\beta_{0}\cap\beta_{1}=\emptyset$. By a modification in the neighbourhood of the intersection point followed by another one in the neighourhood of $\beta_{0}$, we get a curve $f_{1}f_{0}(\beta_{0})$ which is disjoint from $\beta_{0}$. By applying a homeomorphism $f_{2}$ supported in the neighborhood of $\alpha_{0}$, we get $f_{2}f_{1}f_{0}(\beta_{0})$ which additionaly coincides with $\beta_{1}$ near $\alpha_{0}$. Now a modification on the complement of $\alpha_{0}\cup\beta_{0}$ (i.e. a homeomorphism $f_{3}$ of fragmentation norm $1$) can send $f_{2}f_{1}f_{0}(\beta_{0})$ to $\beta_{1}$. The existence of $f_{2}$ and $f_{3}$ is an application of a Schoenflies-like theorem (see for instance [Siebenmann] for an account of the classical case, and [Homma] for a very general case.
@@
We note that this corollary proves the easy direction of the following result due to Emmanuel Militon [Militon]. Here $\mathrm{diam}$ denotes the diameter in the plane.
Theorem 5.4 ([Militon, Theorem 2.4]).
@@
The phenomenon we exhibited here is completely analogous to behavior seen in mapping class groups. Here, Broaddus–Farb–Putman show that the kernel of the Birman exact sequence is exponentially distorted [BFP].
@@ 6. Fine twist numbers and bounded geodesic image theorems
In the beginning of this section and the next one we discuss a general (orientable) surface $S$ of genus $\geq 1$. We will need only the case of the torus, but the tools developed here are likely useful in the general setting. In higher genus, edges of the curve graph corresponds to disjointness, see [Dagger1].
@@ 6.2. Bounded geodesic image theorems
We have the following bounded geodesic image theorems, optimizing either for minimizing the projection bound, or the distance to the projection annulus. Both of these are inspired by Masur-Minsky’s bounded geodesic image theorem, originally proved in [MM2]. Also compare [LongTan] for a bounded geodesic image theorem in higher genus fine curve graphs projecting to non-annular subsurfaces. In the following lemmas, $\mathrm{diam}$ denotes the diameter of sets of vertices in ${\mathcal{A}^{\dagger}}(A)$, and $B_{r}(\alpha)$ denotes a ball in the fine curve graph.
@@
We thank Katie Mann for pointing out to us that this second version follows easily from the first one; our initial proof was an adaptation of Richard Webb’s argument in the usual curve graph ([Webb]).
@@ 7.2. Quasi-paths
Let us first give a more concrete, geometrical characterization of being at almost distance at most 11, which will show in particular that this is a symmetric notion. From now on we fix an orientation on our base curve $\alpha_{0}$, and orient each curve isotopic to $\alpha_{0}$ accordingly, so that the oriented curves are isotopic. Consider the infinite cyclic cover $A=\mathbb{R}\times\mathbb{S}^{1}$ as above where $\alpha_{0}$ lifts to a closed curve, and let $T:A\to A$ denote the covering transformation $(x,y)\mapsto(x+1,y)$. We may choose the orientation so that for each lift $\widetilde{\alpha}$ of $\alpha\in\mathcal{C}^{\dagger}_{0}(T^{2})$, the lift $T\widetilde{\alpha}$ is on the right-hand side of $\widetilde{\alpha}$. Given two essential simple closed curves $a,b$ in $A$, we write $a<b$ if $b$ is included in the open domain bounded by $a$ on its right-hand side, and $a\leq b$ if $b$ is included in the closure of this domain.
@@ 7.4. Operations on curves in the infinite annulus
Denote $\mathcal{C}^{\dagger}(A)$ the set of essential simple closed curve in $A$, which corresponds to curves in the plane that surrounds $\infty$. Given two curves $\alpha,\beta\in\mathcal{C}^{\dagger}(A)$, we denote $\alpha\wedge\beta$ the boundary of $R(\alpha\cup\beta)$. A theorem of Kerekjarto (see [kerekjarto]) asserts that $\alpha\wedge\beta$ is an element of $\mathcal{C}^{\dagger}(A)$, the remarkable fact being that this holds without any regularity nor transversality assumption. Thus we have an operation $\wedge:\mathcal{C}^{\dagger}(A)\times\mathcal{C}^{\dagger}(A)\to\mathcal{C}^{\dagger}(A)$. Note the following caracterization, which is an immediate consequence of the Schoenflies theorem: a point $x$ belongs to $\alpha\wedge\beta$ if and only if (1) it belongs to $\alpha\cup\beta$, and (2) it is accessible from $\infty$ in $R(\alpha)\cap R(\beta)$, that is, there is an embedding with $\rho(0)=x$, $\rho(+\infty)=+\infty$, and $\rho((0,1))\subset R(\alpha)\cap R(\beta)$. We call $\rho$ a witness ray for $x$ in $\alpha\wedge\beta$.
@@ 8. The marking graph estimate
The aim of this section is to prove an upper bound for the distance $d(m,m^{\prime})$ in the marking graph inspired by foundational Masur-Minsky distance formula [MM2] in the “usual” marking graph.
@@
9. Metric WPD property, Mann-Rosendal distance estimates
@@ 9.2. Proof of the metric WPD property
This subsection is dedicated to the proof of the metric WPD property (Theorem 6).
@@
In short, we get the result by showing that $h=g^{-m}f^{n}$ belongs to $Z_{2b}(\alpha,f^{n}\alpha)$, and then applying the WPD. Here are the details. Let $b,f,\alpha$ be as in the statement. By Theorem 6, $f$ satisfies the metric WPD property. Applying this to $\alpha$ and $b$ provides a positive integer $n_{0}$ and a positive number $D$. Let $g\in\mathrm{Homeo}_{0}(S)$, $\beta\in\mathcal{C}^{\dagger}(S)$, $n\geq n_{0}$, $m\in\mathbb{Z}$, and consider the map $h=g^{-m}f^{n}$. By the triangular inequality and the fact that $f$ and $g$ act by isometries, we deduce that $h\in Z_{2b}(\alpha,f^{n}\alpha)$: indeed,
@@ 10.3. Semi-continuity
Compared to the proofs of Theorems 4 and 5, which essentially involved only the two maps $f$ and $g$, the main difficulty here is that we need to control the quality of the quasigeodesic axes for the $g_{k}$’s. This will be achieved via Lemma 3.6.
@@ 10.4. Metric characterization of fixed points of loxodromic
The following statement contains Corollary 1.3 from the introduction (when applied to $f^{-1}$).
@@
Let $D=\widetilde{d}(h,\mathrm{Id})$, and let us fix a lift $\widetilde{\alpha}_{0}$ of $\alpha_{0}$ in the universal cover. From the hypothesis it follows that $h\alpha_{0}$ has a lift which is between $\widetilde{\alpha}_{0}-(0,D)$ and $\widetilde{\alpha}_{0}+(0,D)$. This lifts meets at most $2D+1$ lifts of $\alpha_{0}$. Now the claim follows from [Dagger2, Lemma 4.5]. ∎
@@
11. Stabilizers of points at infinity
In this section we study the stabilizer of the fixed point at infinity of a loxodromic element, proving Corollaries 1.4 and 1.5 and Theorem 1.
Given $f\in\mathrm{Homeo}_{0}(T^{2})$, we define the forward shadowing relation on the torus defined by
$x\sim_{f^{+}}y\Leftrightarrow\text{ the sequence }(\|\widetilde{f}^{n}(\widetilde{x})-\widetilde{f}^{n}(\widetilde{y})\|_{\infty})_{n\geq 0}\text{ is bounded}$
where the norm is the sup norm in $\mathbb{R}^{2}$, $\widetilde{f}$ denotes any lift of $f$, and $\widetilde{x},\widetilde{y}$ any lifts of $x,y$ (the property does not depend on the choices of the lifts). In this situation we also says that the positive orbits of $\widetilde{x}$ and $\widetilde{y}$ shadows each other. The backward shadowing relation is the forward shadowing relation for $f^{-1}$. We first notice that the forward shadowing equivalence relation is highly non trivial.
Lemma 11.1.
Let $f\in\mathrm{Homeo}_{0}(T^{2})$, assume the rotation set of $f$ has interior. Then there are uncountably many forward shadowing equivalence classes.
Proof.
Let us fix a lift $\widetilde{f}$ of $f$. Assume $x$ and $y$ forward shadows each other, and $x$ has rotation vector $\rho=\mathrm{Rot}_{\widetilde{f}}(x)$, that is
$\lim_{n\to+\infty}\frac{1}{n}\left(\widetilde{f}^{n}(\widetilde{x})-\widetilde{x}\right)=\rho.$
Then $y$ also has rotation vector $\rho$. Now the lemma follows immediately from of a result of Llibre and MacKay (item (iv) of Theorem 1 in [LM]), which says in particular that every element in the interior of the rotation set is the rotation vector of some point.
[…]
Proof of Corollary 1.4.
Let $f,g$ be such that $\xi^{-}(f)=\xi^{-}(g)$, we want to prove that their forward shadowing classes coincide. We fix lifts $\widetilde{f},\widetilde{g}$ of $f$ and $g$. By Corollary 1.3, there exists a sequence of positive numbers $(m_{n})_{n\geq 0}$ such that the sequence $(g^{n}f^{-m_{n}})_{n\geq 0}$ is bounded in the group $\mathrm{Homeo}_{0}(T^{2})$. This means that we can find for each $n\geq 0$ some lift $G_{n}$ of $g^{n}$ such that the sequence of sup norms $(\|G_{n}-\widetilde{f}^{m_{n}}\|_{\infty})_{n\geq 0}$ is bounded by some number $D$. Let $\widetilde{x},\widetilde{y}$ whose positive $\widetilde{f}-$orbits shadow each other: the sequence $(\|\widetilde{f}^{m}(\widetilde{x})-\widetilde{f}^{m}(\widetilde{y})\|)_{m\geq 0}$ is bounded by some number $D^{\prime}$. Let us check that the positive $\widetilde{g}-$orbits of $\widetilde{x}$ and $\widetilde{y}$ shadow each other. For each $n\geq 0$ we have $\widetilde{g}^{n}\widetilde{x}-\widetilde{g}^{n}\widetilde{y}=G_{n}\widetilde{x}-G_{n}\widetilde{y}$, and we deduce that this vector is bounded by $D+2D^{\prime}$. ∎
Proof of Corollary 1.5.
Let $f,h,\widetilde{f},\widetilde{h}$ be as in the statement. We first observe that if the numbers $\|\widetilde{f}^{n}-\widetilde{f}^{n}\widetilde{h}\|$$_{\infty}$ are bounded independently of $n\geq 0$ then it follows immediately that for any point $\widetilde{x}$ in the plane, the positive orbit of $\widetilde{h}\widetilde{x}$ shadows that of $\widetilde{x}$: $h$ preserves each forward shadowing class of $f$. This proves the last sentence of the corollary.
We now prove the equivalence between the two properties. First assume the sup norms $\|\widetilde{f}^{n}-\widetilde{f}^{n}\widetilde{h}\|$$_{\infty}$ are bounded independently of $n\geq 0$. Since $\widetilde{h}^{-1}$ is at bounded distance from the identity in the plane, the norms $\|\widetilde{f}^{n}-\widetilde{h}^{-1}\widetilde{f}^{n}\widetilde{h}\|$$_{\infty}$ are also bounded. Then the distances $\widetilde{d}(f^{n},h^{-1}f^{n}h)$ are also bounded, and the sequence $(f^{n}h^{-1}f^{-n}h)_{n\geq 0}$ is bounded in the group $\mathrm{Homeo}_{0}(T^{2})$. By Corollary 1.3 we get $\xi^{-}(f)=\xi^{-}(h^{-1}fh)$, and since $\xi^{-}(h^{-1}fh)=h^{-1}(\xi(f))$ we see that $h$ fixes $\xi^{-}(f)$. This proves that the second condition implies the first one.
[…]
$(\|\widetilde{h}\widetilde{f}^{n}\widetilde{h}^{-1}-\widetilde{f}^{n}\|_{\infty})_{n\geq 0}$
[…]
$\widetilde{h}\widetilde{f}^{m_{n}}\widetilde{h}^{-1}=\widetilde{f}^{n}+\tau_{n}+\varepsilon_{n}\ \ (*).$
We will first show that $m_{n}$ is actually equal to $n$ up to a bounded quantity, and that the sequence $(\tau_{n})$ is bounded. For this, we will make use of the Bounded Deviation theorem of Le Calvez-Tal ([LCT], Theorem D), which says that every displacement vector under $\widetilde{f}^{n}$ is at bounded distance from $n$ times the rotation set of $\widetilde{f}$. Remember that the rotation set is convex, and a conjugacy invariant, and that we have the classical upper bound (Lemma 2.4). Thus we can find $D_{1},D_{2}$ such that the following inequalities hold for every $n\geq 1$:
$)n\mathrm{Rot}(\widetilde{f})\subseteq\mathrm{Conv}(\mathrm{Im}(\widetilde{f}^{n}-\mathrm{Id}))\subseteq N_{D_{1}}(n\mathrm{Rot}(\widetilde{f}$))
$)m_{n}\mathrm{Rot}(\widetilde{f})\subseteq\mathrm{Conv}(\mathrm{Im}(\widetilde{h}\widetilde{f}^{m_{n}}\widetilde{h}^{-1}-\mathrm{Id}))\subseteq N_{D_{2}}(m_{n}\mathrm{Rot}(\widetilde{f}$))
where $N_{D}(X)$ denotes the $D$-neighborhood of a subset $X$ of the plane (remember that the rotation sets are convex and compact, and note that the $D$-neighborhood of a convex set is convex). Plugging in equality (*), we get
$\mathrm{Im}(\widetilde{h}\widetilde{f}^{m_{n}}\widetilde{h}^{-1}-\mathrm{Id})\subseteq N_{D_{0}}(\mathrm{Im}(\widetilde{f}^{n}-\mathrm{Id}))+\tau_{n}$
$\mathrm{Im}(\widetilde{f}^{n}-\mathrm{Id})\subseteq N_{D_{0}}(\mathrm{Im}(\widetilde{h}\widetilde{f}^{m_{n}}\widetilde{h}^{-1}-\mathrm{Id}))-\tau_{n}.$
Combining this with the first inequalities yields
$m_{n}\mathrm{Rot}(\widetilde{f})\subseteq N_{D_{0}+D_{1}}(n\mathrm{Rot}(\widetilde{f}))+\tau_{n}$
$n\mathrm{Rot}(\widetilde{f})\subseteq N_{D_{0}+D_{2}}(m_{n}\mathrm{Rot}(\widetilde{f}))-\tau_{n}.$
(we have used the general equalities $N_{D_{0}}(N_{D_{1}}(Y))=N_{D_{0}+D_{1}}(Y)$ and $\mathrm{Conv}(N_{D}(Y))=N_{D}(\mathrm{Conv}(Y))$). We first deduce than $m_{n}$ is a bounded distance from $n$. Let $\delta$ denotes the diameter of the rotation set of $f$. Since adding $\tau_{n}$ does not change the diameter, we get
$m_{n}\delta\leq n\delta+2(D_{0}+D_{1}),\ \ n\delta\leq m_{n}\delta+2(D_{0}+D_{2})$
and thus
$|m_{n}-n|\leq 2\frac{D_{0}+D_{1}+D_{2}}{\delta}:=D_{3}.$
Now let us prove that $(\tau_{n})$ is bounded. Since $|m_{n}-n|$ is bounded, the sequence $(\|\widetilde{f}^{m_{n}}-\widetilde{f}^{n}\|_{\infty})_{n\geq 0}$ is also bounded. This means that in equality (*) above we may now take $m_{n}=n$, up to increasing the bound on $\varepsilon_{n}$. The last above inclusion now reads
$)n\mathrm{Rot}(\widetilde{f})+\tau_{n}\subseteq N_{D_{0}+D_{2}}(n\mathrm{Rot}(\widetilde{f}$))
which immediately implies that $\|\tau_{n}\|$ is bounded by $D_{0}+D_{2}$.
Finally, up to increasing the bound on $\varepsilon_{n}$ again we may take $\tau_{n}=0$, and equality (*) now reads
$\widetilde{h}\widetilde{f}^{n}\widetilde{h}^{-1}=\widetilde{f}^{n}+\varepsilon_{n}\ \ (*)$
which means that the sequence
$(\|\widetilde{h}\widetilde{f}^{n}\widetilde{h}^{-1}-\widetilde{f}^{n}\|_{\infty})_{n\geq 0}$
is bounded. This completes the proof.
We now prove our Tits alternative in $\mathrm{Homeo}_{0}(T^{2})$. This relies on Corollary 1.5 and on the following classical lemma.
Lemma 11.2.
Let $G$ be a group acting on a Gromov-hyperbolic metric space $X$, and $f\in G$ be a loxodromic element. Then one of the following holds:
  1. (1)
    $G$ fixes $\{\xi^{-}(f),\xi^{+}(f)\}$,
  2. (2)
    $G$ fixes $\xi^{-}(f)$,
  3. (3)
    $G$ fixes $\xi^{+}(f)$,
  4. (4)
    $G$ contains an element $h$ such that the endpoints of the loxodromic element $hfh^{-1}$ are disjoint from those of $f$.
Proof of theorem 1.
Let $G$ be a subgroup of $\mathrm{Homeo}_{0}(T^{2})$ containing a loxodromic element $f$. We explore the four possibilities of Lemma 11.2. If $G$ fixes $\xi^{-}(f)$, then Corollary 1.5 applies, and we get that $G$ leaves invariant each forward shadowing equivalence class. Symmetrically, if $G$ fixes $\xi^{+}(f)$, then $G$ leaves invariant each backward shadowing equivalence class. If $\{\xi^{-}(f),\xi^{+}(f)\}$ is invariant by $G$ then $G$ or an index two subgroup of $G$ fixes both points, and thus preserves each class of both the backward and the forward shadowing relation. In the remain case, by the ping-pong lemma $G$ contains a free group generated by $f^{n}$ and $hf^{n}h^{-1}$, for $h$ given by the lemma and some $n>0$. ∎
Proof of Lemma 11.2.
We denote $\xi^{-}=\xi^{-}(f),\xi^{+}=\xi^{+}(f)$. We assume none of the three first possibilities hold, and we search for some element $h$ fulfilling the fourth condition. By hypothesis, the $G$-orbit of $\xi^{-}$ contains some $x^{-}\not\in\{\xi^{-},\xi^{+}\}$, and symmetrically the orbit of $\xi^{+}$ contains some $x^{+}\not\in\{\xi^{-},\xi^{+}\}$. We want to find some element $h$ of $G$ such that $h\xi^{\pm}\neq\xi^{\pm}$.
Let $g\in G$ be such that $g\xi^{+}=x^{+}\neq\xi^{\pm}$. If $g\xi^{-}\not=\xi^{\pm}$ then we can take $h=g$ and are done. It remains to consider case (1) when $g\xi^{-}=\xi^{+}$ and case (2) when $g\xi^{-}=\xi^{-}$.
First assume we are in case (1). Then $g^{2}\xi^{-}=x^{+}\not\in\{\xi^{-},\xi^{+}\}$. In the case (1.1) where $g^{2}\xi^{+}\neq\xi^{-}$ then also $g^{2}\xi^{+}\neq\xi^{+}$ (since otherwise we would have $x^{+}=g\xi^{+}=g^{-1}\xi^{+}=\xi^{-}$, contradiction). In this case we can take $h=g^{2}$. In the case (1.2) when $g^{2}\xi^{+}=\xi^{-}$ then let $g^{\prime}=fg$. Since $f$ has no fixed point except $\xi^{\pm}$, we get that
$x^{\prime+}:=g^{\prime}\xi^{+}=fx\neq\xi^{\pm},\ \ g^{\prime}\xi^{-}=\xi^{+},\ \$
so we are in case (1) with $g$ replaced by $g^{\prime}$, and now $g^{\prime 2}\xi^{+}\neq\xi^{-}$ (otherwise $g^{\prime}\xi^{+}=g^{\prime-1}\xi^{-}$, that is, $x^{\prime+}=x^{+}$ which is impossible since $f$ does not fix $x^{+}$). So we are in case (1.1) with $g^{\prime}$ replaced by $g$, and we can take $h=g^{\prime 2}$.
Let us consider case (2) when $g\xi^{-}=\xi^{-}$. Then we use the point $x^{-}$ defined above. Let $g^{\prime}\in G$ be such that $g^{\prime}\xi^{-}=x^{-}\neq\xi^{\pm}$. If $g^{\prime}\xi^{+}\neq\xi^{\pm}$ then we can take $h=g^{\prime}$. If $g^{\prime}\xi^{+}=\xi^{-}$ then we are in a case symmetric to the above case (1), and we can take either $h=g^{\prime 2}$ or $h=(f^{-1}g^{\prime})^{2}$. It remains to consider the case when $g^{\prime}\xi^{+}=\xi^{+}$. Remember that we are in the case when $g$ fixes $\xi^{-}$. Let $n>0$ and consider $h=g^{\prime}f^{n}g$. We have $h\xi^{-}=x^{-}\neq\xi^{\pm}$, and $h\xi^{+}=g^{\prime}f^{n}x^{+}$. This last point is not $\xi^{+}$ since $\xi^{+}$ is fixed by $g^{\prime}$ and by $f$. Furthermore, we can assume that $f^{n}x^{+}$ is arbitrarily close to $\xi^{+}$ by taking $n$ arbitrarily large, and then $h\xi^{+}$ is also close to $\xi^{+}$, in particular it is distinct from $\xi^{-}$. So again $h$ suits our needs. This completes the proof. ∎